Suppose you double the radius of a right cylinder. Explain why the surface area is not doubled using the proper formula.
Answer by Theo(13342)


the formula for surface area of a right cylinder is:

sa1 = 2*pi*r*h + 2*pi*r^2

double the surface area and you get:

sa2 = 2*(2*pi*r*h + 2*pi*r^2) which becomes:

sa2 = 4*pi*r*h + 4*pi*r^2

double the radius of the original formula and you get:

sa3 = 2*pi*(2*r)*h + 2*pi*(2*r)^2 which becomes:

sa3 = 2*pi*2*r*h + 2*pi*2^2*r^2 which becomes:

sa3 = 4*pi*r*h + 2*pi*4*r^2 which becomes:

sa3 = 4*pi*r*h + 8*pi*r^2.

if the surface area is doubled when you double the radius, then sa2 would be equal to sa3.

it is not.

sa2 = 4*pi*r*h + 4*pi*r^2
sa3 = 4*pi*r*h + 8*pi*r^2